OpenStudy (anonymous):

A storage charges 32 dollars o print 80 greeting cards and .33 for each additional greeting card Part A Write an equation to find the cost t of 80 greeting cards and c, additional cards Answer Part B Use the equation you wrote in part A to find the total cost of 90 greeting cards and work Answer. Plz plz help seventh grade math

OpenStudy (anonymous):

Let x be the cost of printing cards. If y additional cards are printed Then, x= 32+ 0.33*y Where y is Cards printed above 80 minus 80

OpenStudy (anonymous):

Ok yeah i got the first part

OpenStudy (anonymous):

But how do u do the secound equation can u show it to me

OpenStudy (anonymous):

Would the secound be t= 32+.033(90)

OpenStudy (anonymous):

80 cards cost you 32 dollars. There's been no mention of less. So you assume it to be 32 for even less cards. Additional cards here= 90-80=10 Therefore, t=32+0.33*c Where c=(90-80)

OpenStudy (anonymous):

Ohhhb thanks so much i love u bro

OpenStudy (anonymous):

Umm... That's awkward. But My pleasure anyways